# Disentangled Source-Free Personalization for Facial Expression Recognition with Neutral Target Data

### v1

@@ 3.3 Objective Function
where $y_{\text{g}}$ is the label of the generated image, $p_{\text{id}}$, $p_{\text{exp}}$, and $p_{\text{g}}$ represent the distributions of identity images, expression images, and generated images, respectively and $c_{\text{id}}$, $c_{\text{exp}}$, and $c_{\text{g}}$ are the number of identity, expression, and generated image classes.
@@
$\begin{split}\mathcal{L}_{\text{adv}}=-\lambda_{3}\mathbb{E}_{(\mathbf{I}_{\text{id}},y_{\text{id}})}\sum_{i=1}^{c_{\text{id}}}\mathbbm{1}[y_{\text{id}}=i]\log(D_{\text{id}}(\mathbf{I}_{\text{d}}))\\ -\lambda_{4}\mathbb{E}_{(\mathbf{I}_{\text{exp}},y_{\text{exp}})}\sum_{e=1}^{c_{\text{exp}}}\mathbbm{1}[y_{\text{exp}}=e]\log(D_{\text{exp}}(\mathbf{I}_{\text{g}}))\end{split}$ (6)
@@
$\mathcal{L}_{\text{gen}}=\lambda_{1}\mathcal{L}_{\text{rec}}+\lambda_{2}\mathcal{L}_{\text{per}}+\mathcal{L}_{\text{adv}}$ (7)
In Eqs. (6) and (7), $\lambda_{\text{1}}$ to $\lambda_{\text{4}}$ are hyperparameters used to adjust the impact of individual losses.
Table 1: Accuracy (%) of the proposed method and SOTA methods on the BioVid dataset for 10 target subjects with all 77 sources. Bold text shows the highest accuracy. Note: SFDA (77-shot) indicates that 77 prototypes were used from the source data.
@@
Source only Source model 66.11 55.55 86.36 85.11 87.11 59.00 75.66 70.44 52.66 48.22 68.62
@@
4 Results and Discussion
4.1 Experimental Methodology
We evaluate the performance of our model using the BioVid Heat Pain dataset [46] and the UNBC-McMaster Shoulder Pain dataset [30].
The BioVid dataset is divided into five distinct parts, each varying in terms of subjects, labeling, modalities, and tasks. In our experimentation, we use Part A of the dataset, which includes data from 87 subjects, encompassing four pain levels (PA1, PA2, PA3, and PA4) in addition to a neutral (BL1). The modalities in Part A consist of frontal video recordings and biological signals. In this study, we focus on the highest pain level and the neutral state, using only the frames extracted from frontal video recordings. Each subject has 20 videos per class, each lasting 5.5 seconds. According to [49], the PA4 dataset does not show any facial activity in the first two seconds of the videos and has demonstrated that initial pain intensities do not elicit any facial activities. They suggest focusing solely on ’no pain’ and the highest pain intensities. Therefore, we only consider frames after the first two seconds to exclude the initial part of the sequence that did not indicate any response to pain. In summary, we used 87 subjects with two classes (no pain and pain), 20 videos per class, and 75 frames per video.
Experimental Protocol. In our experiments, we divide both datasets into two domains: source and target. The first experiment is conducted using the BioVid dataset, where the source domain comprises 77 subjects and the target domain includes the remaining 10 subjects, with each subject treated as an individual target domain. To evaluate the performance of our approach, we follow a three-phase methodology: source training, target adaptation, and target testing. During the source training phase, the model is trained using the 77 source subjects, each with both neutral and pain level 4 images. In the adaptation phase, the model is adapted to the target data, consisting of neutral videos of the 10 remaining subjects. Finally, in the target testing phase, the model is tested on each target subject using 12 videos representing both neutral and pain expressions.
In the second experiment, we use the UNBC-McMaster dataset. In this case, 20 subjects are designated as the source domain, while the remaining five subjects form the target domain. Similar to the BioVid experiment, the model is first trained on the source subjects, adapted to the target data, and tested on each target subject individually.
@@ 4.1 Experimental Methodology
Experimental Settings. To validate our approach, several settings are explored: (1) Source only, where a pre-trained model on the source domain is directly tested on the target domain without any adaptation. It serves as a baseline for evaluating how well the model generalizes to unseen data; (2) SFDA, where only neutral images from the target subjects are used for adaptation, while testing is performed on both neutral and pain images. This setting is introduced to investigate how well the model performs when adapted using only neutral images and to compare its results to a scenario where both neutral and pain images are available for adaptation; (3) SFDA ($k$-shot), where we utilized target neutral data and $k$ prototypes (images) from the source domain. The image prototypes from the source data were selected using $k$-means clustering, with $k$ being the number of subjects in the source domain. Therefore, one prototype was selected for each subject. The goal here is to investigate whether pain prototypes, drawn from the source domain, can effectively guide the model in recognizing pain in the target domain; and (4) Oracle, an upper-bound scenario, achieved by fine-tuning the source model with labeled neutral and pain images from the target data.
@@
Table 2: Accuracy (%) of the proposed method and SOTA methods on the UNBC-McMaster dataset for five target subjects with all 19 sources. Bold text shows the highest accuracy. Note: SFDA (19-shot) indicates that 19 prototypes were used from the source data.
@@
Source only Source model 68.56 78.32 60.37 65.54 33.72 61.30
@@
Implementation Details. In our experiments, we used the Inception-V3 CNN with Instance Normalization as the backbone for the encoders. We trained the entire model on the source data using a learning rate of $10^{-5}$ for 100 epochs. To adapt the model to the target subjects, we trained with a learning rate of $10^{-4}$ for 25 epochs, employing a batch size 32. For the expression transfer, we utilized two inputs corresponding to the expression and identity labels. In our experiments, we maintained consistency by using the same dataset for both the identity and expression encoders. For image preprocessing, we resized the input images to $128\times 128$ pixels and applied data augmentation techniques, including random cropping, horizontal flipping, and color adjustments during the training phase. Throughout our experiments, we set the hyperparameters empirically as $\lambda_{1}=5$, $\lambda_{2}=1$, $\lambda_{3}=0.2$, and $\lambda_{4}=0.8$. For $k$-shot SFDA, we set $k=77$ for the BioVid dataset and $k=19$ for the McMaster dataset, which were also determined empirically.
Table 3: Accuracy (%) of the proposed method and state-of-the-art methods on the BioVid dataset for 10 target subjects with all 19 sources from UNBC-McMaster. Bold text shows the highest accuracy.
@@
Source only Source model 50.00 50.00 50.11 50.33 57.89 52.78 58.00 79.44 65.11 48.22 56.48
@@
4.2 Comparison with SOTA Methods
@@
We compared the performance of our DSFDA method with SOTA approaches on the BioVid dataset under various experimental settings, as detailed in Table 1. The Source-only model serves as a lower-bound baseline, where the pre-trained source model is directly tested on target domain data without any adaptation. In the SFDA, the source model is adapted to the target data using only neutral video frames from the target domain. In SFDA ($k$-shot), real neutral images with representative pain prototypes from the source domain are used to address the missing classes in the adaptation phase. The goal is to investigate whether pain prototypes drawn from the source domain can effectively guide the model in recognizing pain in the target domain.
@@
As illustrated in Table 1, the performance of the source only model, which does not utilize any adaptation, is significantly lower compared to the results achieved after adaptation using either synthetic pain images or pain prototypes. This highlights the importance of domain adaptation in addressing the gap between the source and target domains. As expected, the SFDA adaptation setting, which utilizes only short neutral video frames for adaptation, significantly degrades performance, yielding even worse results than the source-only model in existing SOTA methods. When critical expression classes (pain) are missing from the target data, two SOTA methods, SHOT [28], and NRC [54], experience substantial drops in accuracy, with reductions to 18.27% and 8.31%, respectively. These findings highlight the limitations of relying solely on neutral data for adaptation and the challenges models face when key emotional classes are unavailable during adaptation. We used SFDA methods for comparison, which closely align with our proposed method. Our proposed DSFDA method achieves 11.62% higher accuracy than the source-only model due to the generation and adaptation of missing classes in one step through a disentanglement framework. These methods help separate identity-specific and expression-related features, allowing the model to focus on subtle variations in expressions. This leads to better generalization and improved performance, even when pain-specific data is missing. In SFDA ($77$-shot), we introduced source domain prototypes alongside neutral target data for adaptation, selecting representative samples from the source subjects to address the absence of pain data in the target domain. Our proposed DSFDA method outperformed both SHOT and NRC when using these prototypes, demonstrating the effectiveness of leveraging them to bridge the domain gap. Currently, we use a single image prototype for each subject from the source data. The final setting, referred to as the Oracle, involves adapting the source model using both real neutral and pain images from each target subject. This setting serves as the upper bound for adaptation performance, providing the model full access to all relevant class information in the target domain. By utilizing both neutral and pain data during adaptation, the Oracle demonstrates the model’s maximum potential when no data is missing, serving as a benchmark for comparison against more challenging settings where class information is incomplete or synthetic.
@@
The performance of the proposed DSFDA method is compared with SOTA approaches on the UNBC-McMaster shoulder pain dataset under four experimental settings in Table 5. DSFDA consistently outperforms both the source-only baseline and other SOTA methods, achieving an average accuracy of up to 78.75% in the prototype setting. Notably, DSFDA attains 77.70% accuracy in the neutral-only setting, surpassing SHOT and NRC by over 20%.
@@
Next, we evaluated the performance of the proposed DSFDA method and SOTA methods on the BioVid dataset for 10 target subjects using all 19 sources from UNBC-McMaster. As shown in Table 3, the source-only model achieved an average accuracy of 56.48%. Among the SFDA methods, SHOT [28, 42] and NRC [28, 48] achieved average accuracy of 50.86% and 52.30%, respectively. Notably, the proposed DSFDA method obtained the highest average accuracy of 74.01% and outperformed the other methods across most subjects. For more detailed information on the generation and adaptation processes within a two-stage framework, including visual results and the impact of various clustering methods on selecting source image prototypes, please refer to the supplementary material.
4.3 Ablation Studies
Impact of each loss function. This section analyzes the impact of each loss function on the adaptation of the pre-trained source model to the target subjects. We aim to determine how each loss term contributes to the overall target accuracy by evaluating the model’s performance on target data. When only the perceptual and adversarial losses ($\mathcal{L}_{\text{per}}$ and $\mathcal{L}_{\text{adv}}$) are applied, the model achieves 76.28% accuracy on target data, indicating that these losses promote adaptation but lack the reconstruction constraints needed for better performance. When the reconstruction and adversarial losses ($\mathcal{L}_{\text{rec}}$ and $\mathcal{L}_{\text{adv}}$) are used without perceptual loss, there is a slight improvement in target accuracy (0.28%). This suggests that reconstruction helps the model better adapt to the target domain. Applying both reconstruction and perceptual losses without adversarial loss significantly decreases performance on target data (66.00%), highlighting the importance of adversarial training for effective target domain adaptation. When all three losses are combined, the model achieves the best performance, with 81.88% accuracy on target data. The combination of reconstruction, perceptual, and adversarial losses facilitates effective adaptation to the target domain.
@@ 4.3 Ablation Studies
Impact of generation methods on pseudo-labels. We evaluated the reliability of pseudo-labels produced by SHOT and NRC methods when using FExGAN and ImaGINator to generate the missing classes in the target data. ImaGINator consistently produced a higher ratio of reliable pseudo-labels compared to FExGAN. Specifically, the NRC method achieved 75% reliability with FExGAN and improved to 78% with ImaGINator, while the SHOT method resulted in 70% with FExGAN and reached 80% with ImaGINator. Notably, the combination of SHOT with FExGAN yielded the lowest pseudo-label ratio (70%), suggesting that SHOT is less effective when paired with FExGAN compared to ImaGINator. Furthermore, the NRC-FExGAN configuration outperformed our DSFDA-FExGAN method, likely because the generated pain images—often resembling neutral expressions—received reliable pseudo-labels. Since most real or generated pain images do not visibly express pain, assigning pain labels indiscriminately can introduce misleading representations, ultimately diminishing the model’s performance during testing.
Figure 4: t-SNE visualizations of embedding from source to target subjects for sub-1. Different colors represent different classes (green for neutral and orange for pain). (left) the source-only setting, without adaptation; (right) our DSFDA method after adaptation.
4.4 Visualization of Feature Distributions

### v4

@@ III-B Objective Functions
where $y_{\text{g}}$ is the label of the generated image, $p_{\text{id}}$, $p_{\text{exp}}$, and $p_{\text{g}}$ represent the distributions of identity images, expression images, and generated images, respectively, and $c_{\text{id}}$, $c_{\text{exp}}$, and $c_{\text{g}}$ are the number of identity, expression, and generated image classes.
@@
$\begin{split}\mathcal{L}_{\text{adv}}=-\lambda_{\text{adv1}}\mathbb{E}_{(\mathbf{I}_{\text{id}},y_{\text{id}})}\sum_{i=1}^{c_{\text{id}}}\mathbbm{1}[y_{\text{id}}=i]\log(D_{\text{id}}(\mathbf{I}_{\text{d}}))\\ -\lambda_{\text{adv2}}\mathbb{E}_{(\mathbf{I}_{\text{exp}},y_{\text{exp}})}\sum_{e=1}^{c_{\text{exp}}}\mathbbm{1}[y_{\text{exp}}=e]\log(D_{\text{exp}}(\mathbf{I}_{\text{g}}))\end{split}$ (6)
@@
$\mathcal{L}_{\text{gen}}=\lambda_{\text{rec}}\mathcal{L}_{\text{rec}}+\lambda_{\text{per}}\mathcal{L}_{\text{per}}+\mathcal{L}_{\text{adv}}$ (7)
In Eqs. (6) and (7), $\lambda_{\text{adv1}}$, $\lambda_{\text{adv2}}$, $\lambda_{\text{rec}}$, and $\lambda_{\text{per}}$ are hyperparameters used to adjust the impact of individual losses.
IV Results and Discussion
IV-A Experimental Methodology
The BioVid dataset is divided into five distinct parts, each varying in terms of subjects, labeling, modalities, and tasks. In our experimentation, we use Part A of the dataset, which includes data from 87 subjects, encompassing four pain levels (PA1, PA2, PA3, and PA4) in addition to a neutral (BL1). The modalities in Part A consist of frontal video recordings and biological signals. In this study, we focus on the highest pain level and the neutral state, using only the frames extracted from frontal video recordings. Each subject has 20 videos per class, each lasting 5.5 seconds. According to [60], the PA4 dataset does not show any facial activity in the first two seconds of the videos and has demonstrated that initial pain intensities do not elicit any facial activities. They suggest focusing solely on ’no pain’ and the highest pain intensities. Therefore, we only consider frames after the first two seconds to exclude the initial part of the sequence that did not indicate any response to pain.
The StressID dataset [62] comprises facial video recordings from 54 participants, resulting in approximately 918 minutes of annotated visual data. In this work, we focus solely on the visual modality. Each frame is labeled as either neutral (0) or stress (1), based on the participant’s self-reported stress level. Specifically, labels are assigned using a binary scheme where recordings with a stress score below 5 are marked as neutral, and those with a score of 5 or higher are marked as stressed.
Experimental Protocol. In our experiments, each dataset is divided into two domains: source and target. The first experiment is conducted on the BioVid dataset, where 77 subjects form the source domain and the remaining 10 subjects constitute the target domain, with each target subject treated independently. The evaluation follows a three-phase protocol: source training, target adaptation, and target testing. During the source training phase, the model is trained on data from the 77 source subjects, using both neutral and pain level 4 images. In the adaptation phase, the model is adapted to the target data, which consists of neutral videos from the 10 target subjects. Finally, in the target testing phase, the model is evaluated using 12 videos per target subject, covering both neutral and pain expressions. In the second experiment, we use the UNBC-McMaster dataset, designating 19 subjects as the source domain and the remaining 5 as the target. A similar protocol is applied. The third experiment uses the StressID dataset, where 44 subjects are assigned to the source domain and 10 subjects to the target domain. To further evaluate the generalization capabilities of our method, we conduct a cross-dataset experiment, where the source domain comprises 20 subjects from the UNBC-McMaster dataset, and the target domain includes 10 subjects from BioVid.
TABLE I: Accuracy (%) of the proposed method and SOTA methods on the BioVid dataset for 10 target subjects with all 77 sources. Bold text shows the highest accuracy. Note: SFDA (77-shot) indicates that 77 prototypes were used from the source data.
@@ IV-A Experimental Methodology
Source-only Source model 66.11 55.55 86.36 85.11 87.11 59.00 75.66 70.44 52.66 48.22 68.62
@@
Experimental Settings. To validate our approach, several settings are explored: (1) Source-only, where a pre-trained model on the source domain is directly tested on the target domain without any adaptation. It serves as a baseline for evaluating how well the model generalizes to unseen data; (2) SFDA, where only neutral images from the target subjects are used for adaptation, while testing is performed on both neutral and pain images. This setting is introduced to investigate how well the model performs when adapted using only neutral images and to compare its results to a scenario where both neutral and pain images are available for adaptation; (3) SFDA ($k$-shot), where we utilized target neutral data and $k$ prototypes (images) from the source domain. The image prototypes from the source data were selected using $K$-means clustering, with $k$ being the number of subjects in the source domain. Therefore, one prototype was selected for each subject. The goal here is to investigate whether pain prototypes, drawn from the source domain, can effectively guide the model in recognizing pain in the target domain; and (4) Oracle, an upper-bound scenario, achieved by fine-tuning the source model with labeled neutral and pain images from the target data.
Implementation Details. In our experiments, we used the Inception-V3 CNN with Instance Normalization as the backbone for the encoders. We trained the entire model on the source data using a learning rate of $10^{-5}$ for 100 epochs. To adapt the model to the target subjects, we trained with a learning rate of $10^{-4}$ for 25 epochs, employing a batch size of 32. For the expression transfer, we utilized two inputs corresponding to the expression and identity labels. In our experiments, we maintained consistency by using the same dataset for both the identity and expression encoders. For image preprocessing, we resized the input images to $128\times 128$ pixels and applied data augmentation techniques, including random cropping, horizontal flipping, and color adjustments during the training phase. Throughout our experiments, we set the hyperparameters empirically as $\lambda_{\text{rec}}=5$, $\lambda_{\text{per}}=1$, $\lambda_{\text{adv1}}=0.2$, and $\lambda_{\text{adv2}}=0.8$. For the $k$-shot SFDA setting, we define $k$ as the number of source subjects used during training. Specifically, we set $k=77$ for the BioVid dataset, $k=19$ for the UNBC-McMaster dataset, and $k=44$ for the StressID dataset. All experiments were conducted on a single NVIDIA A100 GPU with 48 GB of memory. The source model was trained for 100 epochs, taking approximately 18 hours in total. Compared to baseline methods, our approach is computationally more efficient, as it performs generation and adaptation in a single step—eliminating the need to generate images first and then use them separately for adaptation.
@@
TABLE II: Accuracy (%) of the proposed method and SOTA methods on the UNBC-McMaster dataset for five target subjects with all 19 sources. Bold text shows the highest accuracy. Note: SFDA (19-shot) indicates that 19 prototypes were used from the source data.
@@
Source-only Source model 68.56 78.32 60.37 65.54 33.72 61.30
@@
TABLE III: Accuracy (%) of the proposed method on the StressID dataset for 10 target subjects with all 44 source subjects. Bold text shows the highest accuracy.
Method (Source→Target) Sub-1 Sub-2 Sub-3 Sub-4 Sub-5 Sub-6 Sub-7 Sub-8 Sub-9 Sub-10 Avg.
Source-only 38.96 41.21 65.53 42.04 55.16 65.51 69.43 60.78 53.62 55.63 54.79
DSFDA 73.47 69.39 87.12 69.74 79.87 87.39 82.80 83.89 75.03 77.39 78.61
DSFDA (44-shot) 75.17 71.09 88.82 71.44 81.57 89.09 84.50 85.59 72.73 79.09 79.91
TABLE IV: Accuracy (%) of the proposed method and state-of-the-art methods on the BioVid dataset for 10 target subjects with all 19 sources from UNBC-McMaster. Bold text shows the highest accuracy.
@@
Source-only Source model 50.00 50.00 50.11 50.33 57.89 52.78 58.00 79.44 65.11 48.22 56.48
@@
IV-B Comparison with SOTA Methods
@@
We compared the performance of our DSFDA method with SOTA approaches on the BioVid dataset under various experimental settings, as detailed in Table I. The Source-only model serves as a lower-bound baseline, where the pre-trained source model is directly tested on target domain data without any adaptation. In the SFDA, the source model is adapted to the target data using only neutral video frames from the target domain. In SFDA ($k$-shot), real neutral images with representative pain prototypes from the source domain are used to address the missing classes in the adaptation phase. The goal is to investigate whether pain prototypes drawn from the source domain can effectively guide the model in recognizing pain in the target domain.
@@
As illustrated in Table I, the performance of the source-only model, which does not utilize any adaptation, is significantly lower than the results achieved after adaptation using either synthetic pain images or pain prototypes. This highlights the importance of domain adaptation in addressing the gap between the source and target domains. As expected, the SFDA adaptation setting, which utilizes only short neutral video frames for adaptation, significantly degrades performance, yielding even worse results than the source-only model in existing SOTA methods. When critical expression classes (pain) are missing from the target data, two SOTA methods, SHOT [18], and NRC [19], experience substantial drops in accuracy, with reductions to 18.27% and 8.31%, respectively. These findings highlight the limitations of relying solely on neutral data for adaptation and the challenges models face when key emotional classes are unavailable during adaptation. We used SFDA methods for comparison, which closely align with our proposed method. Our proposed DSFDA method achieves 11.62% higher accuracy than the source-only model due to the generation and adaptation of missing classes in one step through a disentanglement framework. These methods help separate identity-specific and expression-related features, allowing the model to focus on subtle variations in expressions. This leads to better generalization and improved performance, even when pain-specific data is missing. In SFDA ($77$-shot), we introduced source domain prototypes alongside neutral target data for adaptation, selecting representative samples from the source subjects to address the absence of pain data in the target domain. Our proposed DSFDA method outperformed both SHOT and NRC when using these prototypes, demonstrating the effectiveness of leveraging them to bridge the domain gap. Currently, we use a single image prototype for each subject from the source data. The final setting, referred to as the Oracle, involves adapting the source model using both real neutral and pain images from each target subject. This setting serves as the upper bound for adaptation performance, providing the model full access to all relevant class information in the target domain. By utilizing both neutral and pain data during adaptation, the Oracle demonstrates the model’s maximum potential when no data is missing, serving as a benchmark for comparison against more challenging settings where class information is incomplete or synthetic.
@@
The performance of the proposed DSFDA method is compared with SOTA approaches on the UNBC-McMaster and StressID datasets under four experimental settings in Tables VII and III. DSFDA consistently outperforms the source-only baseline in both datasets. On the UNBC-McMaster dataset, our method achieves 78.75% accuracy in the prototype setting. In the neutral-only setting, DSFDA reaches 77.70%, outperforming SHOT and NRC by over 20%.
TABLE V: Impact of loss weighting on target accuracy on BioVid.
$\lambda_{\text{rec}}$ $\lambda_{\text{per}}$ $\lambda_{\text{adv1}}$ $\lambda_{\text{adv2}}$ Accuracy (%)
on target data
0 1 0.2 0.8 76.28
5 0 0.2 0.8 76.50
5 1 0.0 0.0 66.00
5 1 0.2 0.8 81.88
@@
Next, we evaluated the performance of the proposed DSFDA method and SOTA methods on the BioVid dataset for 10 target subjects using all 19 sources from UNBC-McMaster. As shown in Table IV, the source-only model achieved an average accuracy of 56.48%. Among the SFDA methods, SHOT [18, 43] and NRC [18, 44] achieved average accuracy of 50.86% and 52.30%, respectively. Notably, the proposed DSFDA method obtained the highest average accuracy of 74.01% and outperformed the other methods across most subjects. For more detailed information on the generation and adaptation processes within a two-stage framework, including visual results and the impact of various clustering methods on selecting source image prototypes, please refer to the supplementary material.
To better understand the limitations of our method, we conducted a qualitative analysis of failure cases. We observed that the model tends to struggle with subjects exhibiting subtle or ambiguous expressions, particularly for older subjects (e.g., Sub-10 in the BioVid dataset). In some instances, the generated images lacked fine-grained expression details that impacted the adaptation performance.
IV-C Ablation Studies
Impact of hyperparameter values. To assess the individual contributions of each component in the generator’s objective, we perform an ablation study by varying the loss weighting hyperparameters: $\lambda_{\text{rec}}$ (reconstruction loss), $\lambda_{\text{per}}$ (perceptual loss), $\lambda_{\text{adv1}}$ (identity adversarial loss), and $\lambda_{\text{adv2}}$ (expression adversarial loss). Table V presents the results on the target domain. Disabling any of these terms by setting its coefficient to zero leads to a noticeable performance drop, which demonstrates the complementary roles of each objective. The full configuration yields the highest accuracy, validating the effectiveness of our combined loss formulation.
@@ IV-C Ablation Studies
Figure 4: t-SNE visualizations of embeddings from source to target subjects for sub-1. Different colors represent different classes (green for neutral and orange for pain). (left) the source-only setting, without adaptation; (right) our DSFDA method after adaptation.
IV-D Visualization of Feature Distributions